Due to erratic sales of its sole product—a high-capacity battery for laptop computers—PEM, Inc., has been experiencing financial difficulty for some time. The company’s contribution format income statement for the most recent month is given below:

  

Sales (13,200 units × $30 per unit) $ 396,000
Variable expenses 237,600
Contribution margin 158,400
Fixed expenses 176,400
Net operating loss $ (18,000 )

Required:

1. Compute the company’s CM ratio and its break-even point in unit sales and dollar sales.

2. The president believes that a $6,700 increase in the monthly advertising budget, combined with an intensified effort by the sales staff, will result in an $84,000 increase in monthly sales. If the president is right, what will be the increase (decrease) in the company’s monthly net operating income?

3. Refer to the original data. The sales manager is convinced that a 10% reduction in the selling price, combined with an increase of $30,000 in the monthly advertising budget, will double unit sales. If the sales manager is right, what will be the revised net operating income (loss)?

4. Refer to the original data. The Marketing Department thinks that a fancy new package for the laptop computer battery would grow sales. The new package would increase packaging costs by $0.70 per unit. Assuming no other changes, how many units would have to be sold each month to attain a target profit of $4,200?
Answer #1

1.CM Ratio = Contribution Margin/Sales

= 158400/396000

= 40%

Break even point in unit sales = Fixed costs/CM per unit

= 176400/12

= 14,700 units

Break even point Dollar sales = Fixed costs/CM ratio

=176400/40%

= $441,000

2.Increase in Net Operating Income = Increase in Contribution Margin – Increase in cost

= 84000*40% - 6700

= $26,900

3.

Contribution format Income Statement

Sales 26400*27

712800

Less: Variable expenses

475,200

Contribution Margin

237600

Less: Fixed Expenses

206400

Net Operating Income

31,200

4.Target Profit = $4200

Fixed costs = 176400

Target Contribution Margin = $180,600

Units required to be sold = 180600/(30-18-0.7)

= 15,982.30 units